# MambaTab: A Simple Yet Effective Approach for Handling Tabular Data

### v1

@@ Abstract
Tabular data remains ubiquitous across domains despite growing use of images and texts for machine learning. While deep learning models like convolutional neural networks and transformers achieve strong performance on tabular data, they require extensive data preprocessing, tuning, and resources, limiting accessibility and scalability. This work develops an innovative approach based on a structured state-space model (SSM), MambaTab, for tabular data. SSMs have strong capabilities for efficiently extracting effective representations from data with long-range dependencies. MambaTab leverages Mamba, an emerging SSM variant, for end-to-end supervised learning on tables. Compared to state-of-the-art baselines, MambaTab delivers superior performance while requiring significantly fewer parameters and minimal preprocessing, as empirically validated on diverse benchmark datasets. MambaTab’s efficiency, scalability, generalizability, and predictive gains signify it as a lightweight, “out-of-the-box” solution for diverse tabular data with promise for enabling wider practical applications.
1 Introduction
@@
Tabular data remains the predominant data type across industrial, healthcare, academic, and various other domains due to its structured format, despite the recent trend of using images and natural language processing in machine learning. To handle tabular data, numerous strategies have been developed, including machine learning (ML) techniques that use traditional shallow models as well as newer deep learning (DL) architectures. Foundational models such as convolutional neural networks (CNNs) and Transformers Vaswani et al. (2017) have been actively explored and tailored to tabular data modeling in DL, enabling impactful insights and analytics.
While powerful, state-of-the-art deep tabular models typically require a large number of learning parameters, extensive data preprocessing, and hyperparameter tuning. This demands significant computational and human resources, which can impose barriers to developing and deploying these complex models for tabular data, thereby impeding their wider application. Moreover, almost all existing tabular learning methods, except TransTab Wang and Sun (2022), operate under vanilla supervised learning, requiring identical train and test table structures. They are not well-suited for feature incremental learning where features are sequentially added; under such a setting, they have to either drop new features or old data, leading to insufficient use of training data. It is desirable to have the ability to continuously learn from new features.
@@
To address these challenges, we introduce a new approach for tabular data based on structured state-space models (SSMs) Gu et al. (2021b) Gu et al. (2021a) Fu et al. (2022). These models can be interpreted as a combination of CNNs and recursive neural networks, having advantages of both types of models. They offer parameter efficiency, scalability, and strong capabilities for learning representations from varied data, particularly for sequential data with long-range dependencies. To tap into these potential advantages, we leverage SSMs as an alternative to CNNs or Transformers for modeling tabular data.
@@
Specifically, we leverage Mamba Gu and Dao (2023), an emerging SSM variant, as a critical building block to build a new supervised model for tabular data called MambaTab. This proposed model has several key advantages over existing models. Thanks to Mamba’s innovative approach as an SSM, MambaTab not only requires significantly fewer model weights and exhibits linear parameter growth, but also inherently aligns well with feature incremental learning. Additionally, MambaTab has a simple architecture needing minimal data preprocessing or manual tuning. Finally, MambaTab outperforms state-of-the-art baselines, including Transformers and CNN-based models as well as classic learning models.
@@
We extensively benchmark MambaTab against state-of-the-art tabular data approaches. Experiments under two different settings - vanilla supervised learning and feature incremental learning - on 8 public datasets demonstrate MambaTab’s superior performance. It consistently and significantly outperforms the state-of-the-art baselines, including Transformer-based models, while using a small fraction, typically $\mathbf{<1\%}$, of their parameters.
@@
  • Effective end-to-end training and inference with minimal data wrangling needed, in particular, naturally suitable for feature incremental learning
@@ 2 Related Work
In this section, we briefly review existing approaches for learning from tabular data. We roughly categorize them into three groups based on whether they utilize classical shallow models, deep learning like CNN- or contemporary Transformer-based architectures, or self-supervised learning strategies.
Classic Learning-based Approaches
@@
A variety of models exist based on classic ML techniques such as logistic regression (LR), XGBoost Chen and Guestrin (2016) Zhang et al. (2020), and multilayer perceptron (MLP). For example, an MLP variant called self-normalizing neural network (SNN) Klambauer et al. (2017) uses the scaled exponential linear unit (SELU) specifically for tabular data. SNN neuron activations automatically converge towards zero mean and unit variance, enabling high-level abstract representations.
Deep Learning-based Supervised Models
TabNet Arik and Pfister (2021) is a DL model for tabular data modeling based on attention mechanism. It uses sequential attention to choose which features to attend to at each decision step, enabling interpretability and more efficient learning as the learning capacity is used for the most salient features. TabNet is shown to have high performance on a wide range of non-performance-saturated tabular datasets and yield interpretable feature attributions or insights into the global model behavior. Deep cross networks (DCN) Wang et al. (2017) constructs a new network structure consisting of two parts: a deep network and a cross network. The deep network is a standard feed-forward network that can learn high-order feature interactions. The cross network is a new component that can explicitly apply automatic feature crossing with a special operation called vector-wise cross. DCN is shown to be efficient in learning certain bounded-degree feature interactions.
@@
A variety of models have been developed with Transformers as building blocks. AutoInt Song et al. (2019) uses Transformers to learn the importance of different input features. By relying on self-attention networks, this model can automatically learn high-order feature interactions in a data-driven way. TabTransformer Huang et al. (2020) is also built upon self-attention based Transformers, which transform the embeddings of categorical features into robust contextual embeddings to achieve higher prediction accuracy. The contextual embeddings are shown to be highly robust against both missing and noisy data features and provide better interpretability. Moreover, FT-Transformer Gorishniy et al. (2021) converts categorical features into continuous embeddings using a tokenizer, and models the interactions between the continuous embeddings with Transformers.
Self-Supervised Learning-based Models
@@
Recently several approaches have been developed to pre-train deep learning models using self-supervised strategies. VIME Yoon et al. (2020) introduces a tabular data augmentation method for self- and semi-supervised learning frameworks. It builds a pretext task of estimating mask vectors from corrupted tabular data in addition to the reconstruction pretext task. Transformer is used as a component of the pre-trained model. SCARF Bahri et al. (2021) is a self-supervised contrastive learning technique for pre-training on real-world tabular datasets. It forms views for contrastive learning by corrupting a random subset of features. TransTab Wang and Sun (2022) proposes a novel framework for learning from tabular data across tables with different learning strategies such as self-supervised and feature incremental learning. Transformer is used as an integral component of the framework.
3 Method
@@
[…]
@@
In this section, we present our approach for robust learning of tabular data classification, aiming to improve performance through a simple, efficient, yet effective method. Below we describe each component of our method and the working procedures.
Data preprocessing
@@
We consider a tabular dataset, $\{F_{i},y_{i}\}_{i=1}^{m}$, where the features of the $i$-th sample are represented by $F_{i}=\{v_{i,j}\}_{j=1}^{n}$, its corresponding label is $y_{i}\in\{0,1\}$, and $v_{i,j}$ can be categorical, binary or numerical. We treat both binary and categorical features as categorical and utilize an ordinal encoder for encoding them, as shown in Figure 1. Unlike TransTab Wang and Sun (2022), our method does not require manual identification of feature types such as categorical, numerical, or binary. We keep numerical features unchanged in the dataset and handle missing values by imputing the mode. This preprocessing preserves the feature set cardinality, i.e. $n(F_{i})=n({F_{i}}^{\prime})$), where $n(F_{i})$ and $n({F_{i}}^{\prime})$ are the numbers of features before and after processing. Before feeding data into our model, we normalize values $v_{i,j}\in[0,1]$ using min-max scaling:
${v^{\prime}}_{i,j}=\frac{v_{i,j}-min_{i,j=1}^{i=n,j=m}(v_{i,j})}{max_{i,j=1}^{i=n,j=m}(v_{i,j})-min_{i,j=1}^{i=n,j=m}(v_{i,j})}.$ (1)
Embedding representation learning
[…]
Cascading Mamba Blocks
@@
After getting the normalized embedded representations from layer normalization, we apply ReLU activation Agarap (2018) and pass the resulting values $\{u_{k}^{i}\}$, with $u_{k}^{i}$ being the $k$-th token for example $i$, to a Mamba block Gu and Dao (2023). This maps features $Batch\times Length\times Dimension\rightarrow Batch\times Length\times Dimension$. Here, $Batch$ is the minibatch size; $Length$ refers to the token sequence length, and $Dimension$ is the number of channels for each input token. For simplicity, we use $Dimension=1$ by default and $Length$ matches the output dimension from the embedding learning layer (Figure 1). Although Mamba blocks can repeat ${\mathcal{M}}$ times, we set ${\mathcal{M}}=1$ as our default value. However, we perform sensitivity study for ${\mathcal{M}}=2,\cdots,100$ with stacked Mamba blocks, which are connected with residual connections He et al. (2016), to evaluate their information retention or propagation capacity.
Inside a Mamba block, two fully-connected layers in two branches calculate linear projections ($LP_{1},LP_{2})$. The first branch $LP_{1}$’s output passes through a 1D causal convolution and SiLU activation ${\mathcal{S}}(\cdot)$ Elfwing et al. (2018), then a structured state space model (SSM). The continuous-time SSM is a system of first-order ordinary differential equation, which maps an input function or sequence $u(t)$ to output $y(t)$ through a latent state $h(t)$:
$dh(t)/dt=A\ h(t)+B\ u(t),\quad x(t)=C\ h(t),$ (2)
where $h(t)$ is $N$-dimensional, with $N$ also known as a state expansion factor, $u(t)$ is $D$-dimensional, with $D$ being the $Dimension$ factor or the number of channels, $x(t)$ is usually taken as 1D, and $A$, $B$, and $C$ are coefficient matrices of appropriate sizes. This dynamic system induces a discrete version governing state evolution and SSM outputs given the input token sequence via time sampling at $\{k\Delta\}$ with a $\Delta$ time interval. This discrete SSM version is a difference equation:
$h_{k}=\bar{A}\ h_{k-1}+\bar{B}\ u_{k},\quad x_{k}=C\ h_{k},$ (3)
where $h_{k}$, $u_{k}$, and $x_{k}$ are respectively samples of $h(t)$, $u(t)$, and $x(t)$ at time $k\Delta$, $\bar{A}=\exp(\Delta A)$, and $\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$. For SSMs, diagonal $A$ is often used, and Mamba also makes $B$, $C$, and $\Delta$ linear time-varying functions dependent on the input. In particular, for an input token $u$, $B$ and $C$ are both $Linear_{N}(u)$, and $\Delta$ is $softplus(parameter+Linear_{D}(Linear_{1}(u)))$, with $Linear_{p}(u)$ being a linear projection to a $p$-dimensional space and $softplus$ activation function. With such time-varying coefficient matrices, the resulting SSM possesses context and input selectivity properties Gu and Dao (2023), facilitating Mamba blocks to selectively propagate or forget information along the potentially long input token sequence based on the current token. Subsequently, the SSM output is multiplicatively modulated with $\mathcal{S}(LP_{2})$ before another fully connected projection. As a result, these integrated blocks empower MambaTab for content-dependent feature extraction and reasoning with long-range dependencies and feature interactions.
Output Prediction
@@
In this portion, our method learns representations from the concatenated Mamba blocks’ output $\{x_{k}^{i}\}$ of shape $Batch\times Length\times Dimension$, where $x_{k}^{i}$ is the $k$-th token output for example $i$ in a minibatch. These are projected via a fully connected layer from $Batch\times Length\times Dimension\rightarrow Batch\times 1$, resulting in prediction logit $y^{\prime}_{i}$ for example $i$. With sigmoid activation,
$sigmoid(y_{i}^{\prime})=\frac{1}{1+exp(-y_{i}^{\prime})},$ (4)
we obtain the predicted probability score for calculating AUROC and binary-cross-entropy loss.
4 Experiments
4.1 Datasets, Implementation Details, and Baselines
@@
Datasets To systematically evaluate the effectiveness of our method, we utilize 8 diverse public datasets. We provide the dataset’s details and abbreviations in Table 1. Links to the datasets can be found in Supplementary Table 1. Our default experimental settings follow those of  Wang and Sun (2022). We split all datasets into train (70%), validation (10%), and test (20%) sets.
Implementation Details
@@
To keep the preprocessing simple, we follow the approach described in Section 3, generalizing for all datasets without manual intervention or tuning. Post training-validation, we take the best validation model and use it on the test set for prediction. We set up MambaTab with default hyperparameters and tuned potential hyperparameters for each dataset under vanilla supervised learning. For our default hyperparameters, we set up training for 1000 epochs with early stopping patience = 5. We adopt Adam optimizer Kingma and Ba (2014) and cosine-annealing learning rate scheduler with initial learning rate = $1e^{-4}$. In addition to training hyperparameters, MambarTab also involves other model-related hyperparameters and their default values are: embedded representation size ($Length$) = 32, SSM state expansion factor ($N$) = 32, local convolution width (d_conv) = 4, SSM block expansion factor (${\mathcal{M}}$) = 1.
Baselines
@@
We extensively benchmark our model by comparing against standard and current state-of-the-art methods. These include: LR, XGBoost, MLP, SNN with SELU MLP, TabNet, DCN, AutoInt, TabTransformer, FT-Transformer, VIME, SCARF, and TransTab. More information about them can be found in Section 2. For fair comparison, we follow their architectures and implementation detailed in TransTab Wang and Sun (2022).
Performance Benchmark
@@
With default hyperparameters under vanilla supervised learning, our method denoted by MamabaTab-D achieves better performance than state-of-the-art baselines on many datasets and comparable performance on others with far fewer parameters (Table 4). After tuning hyperparameters, we denote our tuned model by MambaTab-T, whose performance further improves. Moreover, under feature incremental learning, our method substantially outperforms the existing method simply with default hyperparameters. We implement MambaTab in PyTorch and will release code upon acceptance. For evaluation, we use Area Under the Receiver Operating Characteristic (AUROC) following  Gorishniy et al. (2021); Wang and Sun (2022). We obtain probability scores via Equation 4 on model output logits and use them and ground truth labels to calculate AUROC.
Table 1: Publicly available datasets with statistics (positive sample ratio, train, validation (val), test data points) and abbreviations used in this paper.
@@
4.2 Vanilla Supervised Learning Performance
For this setting, we follow the protocols from Wang and Sun (2022) directly using the training-validation sets for model learning and the test set for evaluation. To overcome potential sampling bias, we report average results over 10 runs with different random seeds on each of the 8 datasets. With defaults, MambaTab-D outperforms baselines on 3 public datasets (CG, CA, BL) and has comparable performance to transformer-based baselines on others. For example, MambaTab-D outperforms TransTab Wang and Sun (2022) on 5 out of 8 datasets (CG, CA, DS, CB, BL). After tuning hyperparameters, MambaTab-T achieves even better performance, outperforming all baselines on 6 datasets and achieving the second best on the other 2.
Table 2: Test AUROC results on 8 public datasets for vanilla supervised learning. Results reported here are averaged over 10 runs with random splits for our method. The best achieved results are shown in bold and the second best are shown in underlined.

### v2

# MambaTab: A Plug-and-Play Model for Learning Tabular Data

@@ Abstract
Despite the prevalence of images and texts in machine learning, tabular data remains widely used across various domains. Existing deep learning models, such as convolutional neural networks and transformers, perform well however demand extensive preprocessing and tuning limiting accessibility and scalability. This work introduces an innovative approach based on a structured state-space model (SSM), MambaTab, for tabular data. SSMs have strong capabilities for efficiently extracting effective representations from data with long-range dependencies. MambaTab leverages Mamba, an emerging SSM variant, for end-to-end supervised learning on tables. Compared to state-of-the-art baselines, MambaTab delivers superior performance while requiring significantly fewer parameters, as empirically validated on diverse benchmark datasets. MambaTab’s efficiency, scalability, generalizability, and predictive gains signify it as a lightweight, “plug-and-play” solution for diverse tabular data with promise for enabling wider practical applications.
I Introduction
@@
Tabular data, with its structured format, is widely used in industrial, healthcare, academic, and other domains, despite the rise of image and Natural Language Processing (NLP) techniques in machine learning (ML). Numerous ML strategies, including traditional models and newer deep learning (DL) architectures like multi-layer perceptron (MLP), convolutional neural networks (CNNs), and Transformers [1], have been adapted for tabular data, providing valuable insights and analytics.
State-of-the-art deep tabular models either have limited performance or require many parameters, extensive preprocessing, and tuning, using substantial resources, thus limiting their deployment [2]. Furthermore, most methods require consistent table structures for training and testing and struggle with feature incremental learning, which involves sequentially adding features. This typically necessitates dropping either new features or old data, leading to an underutilization of available information [2]. A model capable of continuous learning from new features is needed.
@@
To address these challenges, we propose a new approach for tabular data based on structured state-space models (SSMs) [3, 4, 5]. These models can be interpreted as a combination of CNNs and recursive neural networks, having advantages of both types of models. They offer parameter efficiency, scalability, and strong capabilities for learning representations from varied data, particularly for sequential data with long-range dependencies. To tap into these potential advantages, we leverage SSMs as an alternative to CNNs or Transformers for modeling tabular data.
Specifically, we leverage Mamba [6], an emerging SSM variant, as a critical building block to build a novel model called MambaTab.
@@
This proposed model has several key advantages over existing models: It not only requires significantly fewer model weights and exhibits linear parameter growth but also inherently aligns well with feature incremental learning. Additionally, MambaTab has a simple architecture that needs minimal data preprocessing. Finally, MambaTab outperforms state-of-the-art baselines, including MLP-, Transformer-, and CNN- based models and classic ML models.
@@
We benchmark MambaTab extensively against leading tabular data models. Experiments under three different settings - vanilla supervised learning, self-supervised learning, and feature incremental learning - on 8 public datasets demonstrate MambaTab’s superior performance. It consistently and significantly outperforms the state-of-the-art baselines, including Transformer-based models, while using a small fraction, typically $\mathbf{<1\%}$, of their parameters.
@@
  • Effective end-to-end training and inference with minimal data wrangling needed.
@@ II Related Work
We review existing tabular data learning approaches, roughly categorizing them into classical models, deep learning (CNN or Transformer-based), and self-supervised strategies.
@@
Classic Learning-based Approaches: A variety of models exist based on classic ML techniques such as logistic regression (LR), XGBoost [7] [8], and MLP. Notably, a self-normalizing neural network (SNN) [9], an MLP variant using the scaled exponential linear unit (SELU), is tailored for tabular data. SNNs stabilize neuron activations at zero mean and unit variance, facilitating high-level abstract representations.
Deep Learning-based Supervised Models: TabNet [10] is a DL model that employs an attention mechanism for tabular data, focusing on the most salient features at each decision step for efficient learning and interpretability. It performs well on various tabular datasets and provides interpretable feature attributions. Deep cross networks (DCN) [11] combine a deep network for learning high-order feature interactions and a cross-network that automatically applies feature crossing via a vector-wise cross operation. DCN efficiently learns bounded-degree feature interactions.
@@
A variety of models have been developed with Transformers as building blocks. AutoInt [12] uses Transformers to learn the importance of different input features. By relying on self-attention networks, this model can automatically learn high-order feature interactions in a data-driven way. TabTransformer [13] is also built upon self-attention based Transformers, which transform the embeddings of categorical features into robust contextual embeddings to achieve higher prediction accuracy. The contextual embeddings are shown to be highly robust against both missing and noisy data features and provide better interpretability. Moreover, FT-Transformer [14] tokenizes categorical features into continuous embeddings and models their interactions using Transformers.
@@
Self-Supervised Learning-based Models Several approaches for pre-training deep learning models using self-supervised strategies have emerged. VIME [15] uses tabular data augmentation for self- and semi-supervised learning, creating pretext tasks to estimate mask vectors from corrupted data and for data reconstruction. SCARF [16] employs a self-supervised contrastive learning technique for tabular datasets, generating views for learning by corrupting random feature subsets. TransTab [2] proposes a novel framework for learning from tabular data across tables with different learning strategies including two pre-training strategies of Vertical-Partition Contrastive Learning (VPCL) via supervised and self-supervised techniques and feature incremental learning. UniTabE [17] pretrains a large network for tabular data using a masked language modeling strategy. Transformer is used as an integral component in both TransTab and UniTabE models.
III Method
III-A Preliminaries
Inside a Mamba block, two fully connected layers in two branches calculate linear projections ($LP_{1},LP_{2})$. The first branch $LP_{1}$’s output passes through a 1D causal convolution and SiLU activation ${\mathcal{S}}(\cdot)$ [18], then a structured state space model (SSM). The continuous-time SSM is a system of first-order ordinary differential equation mapping an input function or sequence $u(t)$ to output $x(t)$ through a latent state $h(t)$:
$dh(t)/dt=A\ h(t)+B\ u(t),\quad x(t)=C\ h(t),$ (1)
where $h(t)$ is $N$-dimensional, with $N$ also known as a state expansion factor, $u(t)$ is $D$-dimensional, with $D$ being the $Dimension$ factor or the number of channels, $x(t)$ is usually taken as $D$ dimensional, and $A$, $B$, and $C$ are coefficient matrices of appropriate sizes. This dynamic system induces a discrete version governing state evolution and SSM outputs given the input token sequence via time sampling at $\{k\Delta\}$ with a $\Delta$ time interval. This discrete SSM version is a difference equation:
$h_{k}=\bar{A}\ h_{k-1}+\bar{B}\ u_{k},\quad x_{k}=C\ h_{k},$ (2)
where $h_{k}$, $u_{k}$, and $x_{k}$ are respectively samples of $h(t)$, $u(t)$, and $x(t)$ at time $k\Delta$, $\bar{A}=\exp(\Delta A)$, and $\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$. For SSMs, diagonal $A$ is often used, and Mamba also makes $B$, $C$, and $\Delta$ linear time-varying functions dependent on the input. With such time-varying coefficient matrices, the resulting SSM possesses context and input selectivity properties [6], facilitating Mamba blocks to selectively propagate or forget information along the potentially long input token sequence based on the current token. Subsequently, the SSM output is multiplicatively modulated with $\mathcal{S}(LP_{2})$ before another fully connected projection.
III-B Architecture of Our Model
@@
Figure 2: Illustration of feature incremental learning setting. Feature Seti, $i=1,2,3$, have incrementally added features. Feature SetX represents the set of features for test data.
@@
In this section, we present our approach for robust learning of tabular data classification, aiming to improve performance through a plug-and-play, efficient, yet effective method. Below we describe each component of our method.
@@
Data preprocessing: We consider a tabular dataset, $\{F_{i},y_{i}\}_{i=1}^{m}$, where the features of the $i$-th sample are represented by $F_{i}=\{v_{i,j}\}_{j=1}^{n}$, its corresponding label is $y_{i}\in\{0,1\}$, and $v_{i,j}$ can be categorical, binary or numerical. We treat both binary and categorical features as categorical and utilize an ordinal encoder for encoding them, as shown in Figure 1. Unlike TransTab [2], our method does not require manual identification of feature types such as categorical, numerical, or binary. Moreover, MambaTab only requires 1 embedding learner module whereas TransTab requires 4 embedding modules. We keep numerical features unchanged in the dataset and handle missing values by imputing the mode. This preprocessing preserves the feature set cardinality, i.e. $n(F_{i})=n({F_{i}}^{\prime})$, where $n(F_{i})$ and $n({F_{i}}^{\prime})$ are the numbers of features before and after processing. Before feeding data into our model, we normalize values $v_{i,j}\in[0,1]$ using min-max scaling.
Embedded representation learning: With processed data, we employ a feed-forward network to learn an embedded representation from the features, providing meaningful inputs to our architecture. Although the ordinal encoder imposes ordered representations for categorical features, not all inherently possess such order. Our embedding learner allows the direct learning of multi-dimensional representations from features without depending on imposed orders. This approach also standardizes input feature dimensions for the downstream Mamba blocks across both training and testing in incremental feature learning scenarios, as shown in Figure 2. While most existing methods, except TransTab, are only capable of learning from a fixed set of features, our method MambaTab can learn and transfer weights from Feature Set1 to Feature Set2, and so on. We also utilize layer normalization [19] instead of batch normalization [20] on the learned embedded representations due to its independence of batch size.
@@
Cascading Mamba Blocks: After getting the normalized embedded representations from layer normalization, we apply ReLU activation [21] and pass the resulting values $\{u_{k}^{i}\}$, with $u_{k}^{i}$ being the $k$-th token for example $i$, to a Mamba block [6]. This map features $Batch\times Length\times Dimension\rightarrow Batch\times Length\times Dimension$. Here, $Batch$ is the minibatch size; $Length$ refers to the token sequence length, and $Dimension$ is the number of channels for each input token. For simplicity, we use $Length=1$ by default and $Dimension$ matches the output dimension from the embedding learning layer (Figure 1). Although Mamba blocks can repeat ${\mathcal{M}}$ times, we set ${\mathcal{M}}=1$ as our default value. However, we perform a sensitivity study for ${\mathcal{M}}=2,\cdots,100$ with stacked Mamba blocks, which are connected with residual connections [22], to evaluate their information retention or propagation capacity. As a result, these integrated blocks empower MambaTab for content-dependent feature extraction and reasoning with long-range dependencies and feature interactions.
@@
Output Prediction: In this portion, our method learns representations from the concatenated Mamba blocks’ output $\{x_{k}^{i}\}$ of shape $Batch\times Length\times Dimension$, where $x_{k}^{i}$ is the $k$-th token output for example $i$ in a minibatch. These are projected via a fully connected layer from $Batch\times Length\times Dimension\rightarrow Batch\times 1$, resulting in prediction logit $y^{\prime}_{i}$ for example $i$. With sigmoid activation, we obtain the predicted probability score for calculating AUROC and binary-cross-entropy loss.
IV Experiments
IV-A Datasets, Implementation Details, and Baselines
@@
Datasets: To systematically evaluate the effectiveness of our method, we utilize 8 diverse public benchmark datasets which are also widely followed by TransTab [2] and UniTabE [17]. We provide the dataset’s details and abbreviations in Table I. Our default experimental settings follow those of  [2]. We split all datasets into train (70%), validation (10%), and test (20%).
@@
Implementation Details: To keep the preprocessing simple, we follow the approach described in Section 2, generalizing for all datasets without manual intervention. For post-training validation, we take the best validation model and use it on the test set for prediction. We set up MambaTab with default hyperparameters and tuned potential hyperparameters for each dataset under vanilla supervised learning. For our default hyperparameters, we set up training for 1000 epochs with early stopping patience = 5. We adopt Adam optimizer [23] and cosine-annealing learning rate scheduler with initial learning rate = $1e^{-4}$. In addition to training hyperparameters, MambaTab also involves other model-related hyperparameters and their default values are: embedded representation size = 32, SSM state expansion factor ($N$) = 32, local convolution width (d_conv) = 4, number of SSM blocks ($\mathcal{M}$) = 1.
@@
Baselines: We extensively benchmark our model by comparing it against standard and current state-of-the-art methods. These include: LR, XGBoost, MLP, SNN with SELU MLP, TabNet, DCN, AutoInt, TabTransformer, FT-Transformer, VIME, SCARF, UniTabE and TransTab. More information about them can be found in Section II. For a fair comparison, we follow their implementation detailed in TransTab [2].
@@
Performance Benchmark: With default hyperparameters under vanilla supervised learning, our method denoted by MambaTab-D achieves better performance than state-of-the-art baselines on many datasets and comparable performance on others with far fewer parameters (Table V). After tuning hyperparameters, we denote our tuned model by MambaTab-T, whose performance further improves. Moreover, under feature incremental learning, our method substantially outperforms the existing method simply with default hyperparameters. We implement MambaTab in PyTorch, which can be found here [^1] https://github.com/Atik-Ahamed/MambaTab. For evaluation, we use Area Under the Receiver Operating Characteristic (AUROC) following  [14, 2, 17].
TABLE I: Publicly available datasets with statistics.
@@
IV-B Vanilla Supervised Learning Performance
For this setting, we follow the protocols from [2] directly using the training-validation sets for model learning-tuning and the test set for evaluation. To overcome potential sampling bias, we report average results over 10 runs with different random seeds on each of the 8 datasets. With defaults, MambaTab-D outperforms baselines on 3 public datasets (CG, CA, BL) and has comparable performance to transformer-based baselines on others. For example, MambaTab-D outperforms TransTab [2] on 5 out of 8 datasets (CG, CA, DS, CB, BL). After tuning hyperparameters, MambaTab-T achieves even better performance, outperforming all baselines on 6 datasets and achieving the second best on the other 2 datasets. In Table II, we did not find performance on the IC dataset from the UniTabE paper. UniTabE-S and UniTabE-FT (Table IV) refer to the cases of training from scratch and utilizing the pre-trained weights and fine-tuning on the target datasets, respectively.
TABLE II: Test AUROC for vanilla supervised learning. The best results are shown in bold and the second best are shown in underlined.